# Nonparametric regression for cost-effectiveness analyses with observational data - a tutorial

### v1

@@ A Simulation experiment
We demonstrate here, through a simple simulation experiment, the consequences of model misspecification when estimating treatment effects. Such experiments are useful because we know the true data-generating process (hence also the true treatment effects) and can hence investigate how a statistical method performs at recovering the truth. R code to replicate the experiment is given at the end.
The goal of our hypothetical analysis is estimating effect of a treatment $t_{i}$ on the healthcare costs $c_{i}$, and there is a single continuous baseline covariate $x_{i}$. Our true model is given by
$c_{i}=t_{i}+3\sin(x_{i})x_{i}^{2}+\epsilon_{i}$ (3)
In line with common practice, our choice of the statistical model will be a simple linear regression:
@@
However, our model in (4) is evidently not correctly specified. The question now arises what happens to our estimates, when we fit model (4) regardless of the misspecification. Do they still provide useful information about the treatment effect?
As mentioned in Section 3, the answer depends on whether the treatment assignment depends on the baseline covariate $x_{i}$. We first assign, to each patient, a probability of $\frac{1}{2}$ of receiving treatment $t_{i}=1$, independently of $x_{i}$. The simulation hence mimics a randomized trial.
We now draw $100$ random datasets from our specified simulation, and then analyse each dataset with the linear regression model. The left panel in Figure A.4 shows a histogram of the estimated treatment effects. Despite the misspecification, the estimates concentrate around the true value $1$.
@@
The probability in (5) is called the propensity score.
The results are shown in the right panel of Figure A.4. Unlike before, we observe a clear bias in the estimates.
While our example was of course rather simplistic, the same patterns can be observed in more complex and realistic simulation experiments. A case in point is the seminal work by Dorie et al. (2019), where a much more elaborate and systematic simulation experiment may be found. The authors simulated data a variety of hypothetical observational studies, using nonlinear data-generating processes. They then had other teams of researchers, who had not seen the true data-generating processes, try to estimate the average treatment effect through a statistical method of their choice. It turned out that all variations of linear regression models performed poorly, as did other kinds of parametric, while some nonparametric regression methods did very well.
To illustrate the utility of nonparametric regression, we rerun the simulation experiment from the previous section, using BART instead of a linear regression. The results are presented in A.5. For each of the two data-generating processes, the inferred treatment effects concentrate around the true value.
@@
1N <- 200
2X <- rnorm(N)
3ATE_linear_sim1 <- rep(NA, N_sim)
4ATE_linear_sim2 <- rep(NA, N_sim)
5ATE_BART_sim1 <- rep(NA, N_sim)
6ATE_BART_sim2 <- rep(NA, N_sim)
7EY_0 <- 3 * sin(X) * X^2
8EY_1 <- EY_0 + 1
9ATE_true <- mean(EY_1 - EY_0)
10
11N_sim <- 100
12test <- matrix(c(rep(X, 2), rep(0, N), rep(1, N)), ncol = 2)
13
14# sim 1: random assignment
15for (i in 1:N_sim){
@@ 43 ATE_BART_sim2[i] <- mean(BART_mod\$y_hat_test_mean[(N+1):(2*N),1] - BART_mod\$y_hat_test_mean[1:N,1])
44}
45
46data_sim <- data.frame(sim = c(rep("randomized",N_sim),rep("non-randomized",N_sim)),
@@ 48 ATE_BART = c(ATE_BART_sim1,ATE_BART_sim2)
49)
50data_sim <- data.frame(sim = c(rep("randomized",N_sim),rep("non-randomized",N_sim)),
@@ 52 ATE_BART = c(ATE_BART_sim1,ATE_BART_sim2)
53)
54data_sim\$sim <- factor(data_sim\$sim, levels = c("randomized", "non-randomized"))
55
56ggplot(data_sim) +
@@
Listing 8: Code for simulation experiment

### v2

@@ A Simulation experiment
We demonstrate here, through a simple simulation experiment, the consequences of model misspecification when estimating treatment effects. Such experiments are useful because we know the true data-generating process (hence also the true treatment effects) and can hence investigate how a statistical method performs at recovering the truth. The R code to replicate the experiment is given at the end.
The goal of our hypothetical analysis is estimating the effect of a treatment $t_{i}$ on the healthcare costs $c_{i}$, and there is a single continuous baseline covariate $x_{i}$. Our true model is given by
$c_{i}=t_{i}+3\sin(x_{i})x_{i}^{2}+\epsilon_{i}.$
In line with common practice, our statistical model of choice is a simple linear regression:
@@
However, our model in Equation (2) is evidently misspecified. This raises the question of what happens to our estimates when we fit model  (2) despite its misspecification. Specifically, we ask whether the estimates can still yield meaningful information about the underlying treatment effect.
As mentioned in Section 3, the answer depends on whether the treatment assignment depends on the baseline covariate $x_{i}$. We first assign to each patient a probability of 0.50.50.5 of receiving treatment $t_{i}=1$, independently of $x_{i}$. The simulation setup hence mimics a randomized trial.
We now draw 100100100 random datasets from the specified simulation and analyse each dataset using the linear regression model. The left panel in Figure A.4 shows a histogram of the estimated treatment effects. Despite the misspecification, the estimates concentrate around the true value 111.
@@
The probability in Equation (3) is called the propensity score (PS).
The results are shown in the right panel of Figure A.4. Unlike before, we now observe a clear bias in the estimates.
While our example was of course rather simplistic, the same patterns can be observed in more complex and realistic simulation experiments. A case in point is the seminal work by Dorie et al. (2019), where a much more elaborate and systematic simulation experiment may be found. The authors simulated data from a variety of hypothetical observational studies, using nonlinear data-generating processes. They then tasked other teams of researchers, who were unaware of the true data-generating processes, to try to estimate the average treatment effect through a statistical method of their choice. It turned out that all variations of linear regression models performed poorly, as did other kinds of parametric models, whereas some nonparametric regression methods performed well.
To illustrate the utility of nonparametric regression, we rerun the simulation experiment from the previous section, using BART instead of a linear regression. The results are presented in A.5. For each of the two data-generating processes, the inferred treatment effects concentrate around the true value, which shows that BART can correctly estimate the treatment effects even when the true relationship is nonlinear.
@@
1rm(list=ls())
2library(subart)
3library(ggplot2)
4
5N_sim <- 100
6N <- 200
7X <- rnorm(N)
8ATE_linear_sim1 <- rep(NA, N_sim)
9ATE_linear_sim2 <- rep(NA, N_sim)
10ATE_BART_sim1 <- rep(NA, N_sim)
11ATE_BART_sim2 <- rep(NA, N_sim)
12EY_0 <- 3 * sin(X) * X^2
13EY_1 <- EY_0 + 1
14ATE_true <- mean(EY_1 - EY_0)
15
16test <- matrix(c(rep(X, 2), rep(0, N), rep(1, N)), ncol = 2)
17test <- as.data.frame(test)
18
19# sim 1: random assignment
20for (i in 1:N_sim){
@@ 48 ATE_BART_sim2[i] <- mean(BART_mod\$y_hat_test_mean[(N+1):(2*N),1] - BART_mod\$y_hat_test_mean[1:N,1])
49}
50
51data_sim <- data.frame(sim = c(rep("randomized",N_sim),rep("non-randomized",N_sim)),
@@ 53 ATE_BART = c(ATE_BART_sim1,ATE_BART_sim2)
54)
55data_sim <- data.frame(sim = c(rep("randomized",N_sim),rep("non-randomized",N_sim)),
@@ 57 ATE_BART = c(ATE_BART_sim1,ATE_BART_sim2)
58)
59data_sim\$sim <- factor(data_sim\$sim, levels = c("randomized", "non-randomized"))
60
61ggplot(data_sim) +
@@
Listing 8: Code for the simulation experiment